# Gradient Boosting Decision Trees on Medical Diagnosis over Tabular Data

### v3

@@ I Introduction
Following the recent application fields, one of the most important environments that utilize tabular data is the medical field, especially several diagnosis tasks using patient records. Here, each patient record can be treated as row entries, and columns have the corresponding attributes of the patients. Possibility of having sparsity and less correlation between each record, robust GBDT methods have gained high popularity in the field [5], [6].
@@ II-A Traditional Machine Learning Models
Various studies have explored the application of ML algorithms in medical diagnosis over tabular data. For example, KNN algorithms are used in many applications due to their simplicity, especially in medical data classification, where the data consist of many samples of medical health [10], when detecting breast cancer [7], and providing diagnosis on diabetes by using a fine-tuned KNN classifier [11]. Another example can be logistic regression, that has been widely used in both continuous and categorical data classifications, including the medical diagnosis field [12], specifically in COVID-19 diagnosis [13], and on many other major chronic disease diagnoses [14]. On the other hand, SVMs are relatively famous for performing medical diagnosis [15], especially when combined with other different ML methods, including several deep neural networks [16], [17]. Moreover, decision trees are one of the most widely used classifiers over tabular data in general [18], and for medical diagnosis specifically, such as coronary artery disease classification [19], chronic kidney disease diagnosis [8]. Their performance can be increased when combined with other deep neural network models, such as training the decision trees using convolutional neural networks (CNN) for COVID-19 diagnosis [20].
@@ IV-A Datasets
We evaluated the performance of GBDT models over 7 medical diagnosis datasets. Dataset information that includes the number of samples, features, classes, and the type of the task is provided in Table I. Cardiovascular Disease Dataset (CD) [53] contains heart disease data that consist of 70000 patient records with 11 independent features, including a binary label. The aim is to build a predictive algorithm that detects the presence of an early-stage heart disease. Heart Failure [54] is another heart-related dataset that contains 299 medical records of different patients who had heart failure, where each of the patients has 13 clinical features. The aim is to predict the death event for the given patient records, i.e., whether there will occur heart failure with respect to the given patient features. Parkinson’s [55] dataset comprises several biomedical voice measurements from a total of 31 patients, where 23 of them have Parkinson’s disease. There is a total of 195 different samples from these patients, where each of them includes 22 features. These features include continuous measurements. It is used to predict the status of a given patient data, which corresponds to being healthy or having Parkinson’s disease. EEG Eye State [56] dataset consists of 14 features for each patient that hold values regarding the EEG measurements. These features include continuous measurements. There are nearly 15000 samples and each of them has a label indicating the eye-closed or eye-open states. It is aimed to predict the state of a patient sample using these EEG measurement features. Eye Movements [57] contains 27 features in which 22 of them comprises features that are commonly used in psychological studies on eye movements. The rest of the features include extra information about the samples, that each of the samples is an assignment. These assignments consist of a question, that is followed up by several answer sentences. The task is to predict the multi-class relevance label of an answer for each assignment, in which the labels are irrelevant, relevant, and the correct answer. Arcene [58] dataset includes several features that indicate the abundance of proteins in human sera, having a given mass value. There are 200 patient samples and nearly 10000 features in total. These features are continuous. The task is to predict patients who have cancer and patients that are healthy, according to the given features. Lastly, Prostate [59] dataset includes continuous features that give information regarding prostate cancer. There are 102 patient samples and nearly 12600 total features. The aim is to predict whether a given patient has prostate cancer or not.
@@ V Discussion
Observing Fig.’s 1 and 2, DL models tend to have higher average training times compared to other ML models due to the complex structures that consist of large number of parameters that needs to be optimized. Specifically, in Fig.’s 1(a) and 2(a), DL models have higher average training times than traditional ML and GBDT models. In Fig.’s 1(b) and 2(b), the above trend continues except CatBoost having increased training time in Fig. 1(b) and higher training time than DL models in Fig. 2(b). The reason behind is that since CatBoost relies on building an ensemble of trees sequentially, due to the existence of high dimensions and intricate patterns in medical datasets, such ensemble models might require deeper and larger number of trees to capture the dependencies in the data, leading to longer training times. Having a relatively higher number of dimensions in the Prostate dataset, CatBoost required a higher average training time than normal, which caused some of the DL models, such as TabNet and VIME, have shorter average training times. In general, GBDT methods employ high overall performance in all of the datasets, yet have a good computational performance by having relatively less average training times.
@@ VI Conclusion
In this work, we investigated the overall superiority of ensemble models, especially GBDTs, over other state-of-the-art tabular DL models and traditional ML methods in medical diagnosis, for several benchmark tabular medical datasets. In our analysis, we explored the trade-offs between performance, computational cost, and ease of optimization between the models. Overall, GBDT models exhibited superior performance compared to traditional ML and other deep architectures. Additionally, GBDTs are computationally convenient due to the less complex architectures compared to the deep neural networks that are considerably complex. Due to the less complex structure, GBDTs were much easier to optimize in comparison. Consequently, GBDT models can be safely used in various medical tasks for providing fast and accurate results in any type of diagnosis. In practice, our findings can facilitate clinical decisions. These insights can help data scientists and medical professionals to optimize model selection based on performance and time efficiency. Within this approach, patient care and delivery of healthcare can be improved since our study provides a comprehensive evaluation across a range of ML and DL models as a representative of real-world clinical challenges.
Despite the significant progress of gradient boosting methods over tabular data, further research can be conducted on benchmarking the advantages of GBDTs across alternative settings in healthcare. These alternative settings have the potential to provide further opportunities for analyzing, utilizing, and implementing the GBDT approach. Moreover, they pretty much compose the field of tabular data analysis due to the suffering of deep architectures, that can be extended to any type of task utilized from using tabular data. Due to the significance of tabular data both to the industry and academia, our findings can provide remarkable assistance in the field.

### v4

@@ I Introduction
Following the recent application fields, one of the most important environments that utilize tabular data is the medical field, especially several diagnosis tasks using patient records. Here, each patient record can be treated as row entries, and columns have the corresponding attributes of the patients. Such medical datasets inherently contain weakly correlated features, meaning that variables often do not exhibit closely associated fluctuations. Understanding these correlations is essential for model interpretation, especially as models grow more complex with numerous interacting variables. Several ML models are interpretable-by-nature such as logistic regression, KNN, and decision trees [5], [6]. In particular, GBDTs provide such interpretable model structures that are crucial in medical diagnosis in terms of providing model transparency and user trust, in contrast to the “black-box” DL models in which the decision-making process is non-intuitive and difficult to understand [7]. Thus, in the possibility of having sparsity and less correlation between each record, robust GBDT methods have gained high popularity in the field [8], [9].
@@ II-A Traditional Machine Learning Models
Various studies have explored the application of ML algorithms in medical diagnosis over tabular data. For example, KNN algorithms are used in many applications due to their simplicity, especially in medical data classification, where the data consist of many samples of medical health [13], when detecting breast cancer [10], and providing diagnosis on diabetes by using a fine-tuned KNN classifier [14]. Another example can be logistic regression, that has been widely used in both continuous and categorical data classifications, including the medical diagnosis field [15], specifically in COVID-19 diagnosis [16], and on many other major chronic disease diagnoses [17]. On the other hand, SVMs are relatively famous for performing medical diagnosis [18], especially when combined with other different ML methods, including several deep neural networks [19], [20]. Moreover, decision trees are widely used classifiers for tabular data due to their ease of interpretation. Their hierarchical structure allows for straightforward visualization and understanding of the decision-making process [21]. For medical diagnosis, decision trees are used for coronary artery disease classification [22] and chronic kidney disease diagnosis [11]. Their performance can be increased when combined with other deep neural network models, such as training the decision trees using convolutional neural networks (CNN) for COVID-19 diagnosis [23].
@@ IV-A Datasets
We evaluated the performance of GBDT models across seven medical diagnosis datasets. Table I provides details such as the number of samples, features, classes, and task types. The Cardiovascular Disease (CD) dataset [56] includes 70,000 patient records with 11 features and a binary label, aiming to detect early-stage heart disease. The Heart Failure dataset [57] contains 299 patient records with 13 clinical features, used to predict the likelihood of a death event due to heart failure. The Parkinson’s dataset [58] includes 195 samples from 31 patients (23 with Parkinson’s), each with 22 biomedical voice measurements, for binary classification. The EEG Eye State dataset [59] has 14 EEG-related features and 15,000 samples, labeled as either eye-open or eye-closed. Eye Movements [60] consists of 27 features (22 used in psychological studies) for classifying answer relevance (irrelevant, relevant, or correct) in multi-class settings. The Arcene dataset [61] contains 200 samples with about 10,000 continuous features related to protein abundance, for binary classification of cancer presence. Lastly, the Prostate dataset [62] includes 102 samples with roughly 12,600 continuous features to predict prostate cancer status.
@@ IV-D Hyperparameter Optimization
We trained the GPU-supported ML and DL models on NVIDIA A100 Tensor Core GPU, where the average training over the folds lasted 342 minutes at most across all datasets.
@@ V Discussion
Observing Fig.’s 1 and 2, DL models tend to have higher average training times compared to other ML models due to the complex structures that consist of large number of parameters that needs to be optimized. Specifically, in Fig.’s 1(a) and 2(a), DL models have higher average training times than traditional ML and GBDT models. Although GBDTs are usually computationally-efficient compared to the DL models, they face challenges in the tradeoff between accuracy and efficiency, mostly in the high feature dimension cases, where the computational complexities become proportional to the number of features [53]. As an example, in Fig.’s 1(b) and 2(b), the low-computation trend of GBDTs continues except CatBoost that has increased training time in Fig. 1(b) and higher training time than DL models in Fig. 2(b). The reason behind is that since CatBoost relies on building an ensemble of trees sequentially, due to the existence of high dimensions and intricate patterns in medical datasets, such ensemble models might require deeper and larger number of trees to capture the dependencies in the data, leading to longer training times. Having a relatively higher number of dimensions in the Prostate dataset, CatBoost required a higher average training time than normal, which caused some of the DL models, such as TabNet and VIME, have shorter average training times. In general, GBDT methods employ high overall performance in all of the datasets, yet have a good computational performance by having relatively less average training times.
@@ VI Conclusion
In this study, we investigated the overall superiority of ensemble models, especially GBDTs, over other state-of-the-art tabular DL models and traditional ML methods in medical diagnosis, for several benchmark tabular medical datasets. In our analysis, we explored the trade-offs between performance, computational cost, and ease of optimization between the models. Overall, GBDT models exhibited superior performance compared to traditional ML and other deep architectures. Additionally, GBDTs are computationally convenient due to the less complex architectures compared to the deep neural networks that are considerably complex. Due to the less complex structure, GBDTs were much easier to optimize in comparison. Consequently, GBDT models can be safely used in various medical tasks for providing fast and accurate results in any type of diagnosis. In practice, our findings can facilitate clinical decisions. These insights can help data scientists and medical professionals to optimize model selection based on performance and time efficiency. Within this approach, patient care and healthcare delivery can be improved since our study provides a comprehensive evaluation across a range of ML and DL models as a representative of real-world clinical challenges.
Despite the significant progress of gradient boosting methods over tabular data, further research can be conducted on benchmarking the advantages of GBDTs across alternative real-life settings in healthcare, as publicly available medical datasets come with limitations such as potential data quality concerns and the static nature of pre-collected data. These alternative settings have the potential to provide further opportunities for analyzing, utilizing, and implementing the GBDT approach. Moreover, they pretty much compose the field of tabular data analysis due to the suffering of deep architectures, that can be extended to any type of task utilized from using tabular data. Due to the significance of tabular data both to the industry and academia, our findings can provide remarkable assistance in the field.